# PerfTracker: Online Performance Diagnosis for Large Model Training

### v3

@@ Abstract.
Troubleshooting performance problems of large model training (LMT) is immensely challenging, due to unprecedented scales of modern GPU clusters, the complexity of software-hardware interactions, and the data intensity of the training process. Existing troubleshooting approaches designed for traditional distributed systems or datacenter networks fall short and can hardly apply to real-world training systems. In this paper, we present PerfTracker, the first online troubleshooting system utilizing fine-grained profiling, to diagnose performance issues of large-scale model training in production. PerfTracker can diagnose performance issues rooted in both hardware (e.g., GPUs and their interconnects) and software (e.g., Python functions and GPU operations). It scales to LMT on modern GPU clusters. PerfTracker effectively summarizes runtime behavior patterns of fine-grained LMT functions via online profiling, and leverages differential observability to localize the root cause with minimal production impact. PerfTracker has been deployed as a production service for large-scale GPU clusters of $O(10,000)$ GPUs (product homepage https://help.aliyun.com/zh/pai/user-guide/perftracker-online-performance-analysis-diagnostic-tool). It has been used to diagnose a variety of difficult performance issues.
copyright: nonedoi: isbn: price:
1. Introduction
@@
Troubleshooting performance problems of Large Model Training (LMT) has become a grand challenge driven by unprecedented increases in AI model sizes and GPU-based training infrastructures. LMT undertakes complex hardware-software interactions; unexpected behavior of hardware components (e.g., GPUs and their interconnects) to software code and configurations (e.g., data loaders, communication, and PyTorch), could lead to performance issues. With rapid scaling of LMT, occurrences of performance problems inevitably increase accordingly, resulting in significant wastage of compute resources and production incidents (Jiang et al., 2024; Qian et al., 2024; Xiong et al., 2024).
Unfortunately, traditional troubleshooting techniques, designed for distributed systems (Zhao et al., 2014; Barham et al., 2004; Terai et al., 2014; Harsh et al., 2023) and data center networks (Liu et al., 2024; Tan et al., 2019; Arzani et al., 2018; Roy et al., 2017), fall short and can hardly help today’s model training systems. First, all existing approaches focus on CPU-based systems and do not work with GPU-centric, heterogeneous systems. Modern LMT systems are orders of magnitude larger in scale than traditional distributed systems and networks in terms of parallel computation. The scale and parallelism impair existing approaches that rely on fine-grained traces and profiling data. For example, an LMT of 10,000 GPUs typically makes $4\times 10^{9}$ function executions per second, which produces terabytes of profiling data. Active measurement (e.g., probing system components (Liu et al., 2024; Tan et al., 2019)) is not an option due to the interference with production workloads. Moreover, most existing approaches focus only on software defects or only on hardware infrastructures; they cannot reason about their complex interactions.
In practice, troubleshooting performance issues of production LMT relies on data from either online monitoring or offline profiling. Unfortunately, neither is ideal and engineers often face difficult trade-offs between observability and cost. Online monitoring has to be coarse-grained to avoid production impact, while offline profiling loses timeliness and requires a reproduction of the observed problem in a small test environment (which is not always possible).
@@
Most online monitors use vendor tools to collect performance counters of hardware components (e.g., GPUs (dcg, 2024), CPU (pcm, 2024), CPU-GPU (dyn, 2024), and NICs (pcm, 2024; mst, 2024)) during LMT. To avoid production impacts, the sampling can only be coarse-grained, typically in seconds. Similarly, monitors on function executions can only record coarse-grained runtime information like forward phase, backward phase, and CUDA kernels (meg, 2019; Jiang et al., 2024). However, coarse-grained observability cannot capture fine-grained operations in LMT, which is critical for performance analysis. In our experience, online monitors can only detect coarse-grained symptoms (e.g., performance degradation in forward/backward phases or in communication modules), but cannot pinpoint fine-grained root causes (e.g., which function or network link caused the degradation).
@@
Offline profiling (e.g., Torch Profiler (tor, 2024), Nsight System (nsy, 2024), and Nsight Compute (nsi, 2024)) records fine-grained operations during LMT execution, including both function executions (e.g., Python call stacks, CUDA executions) and high-frequency hardware sampling (e.g., at 200k Hz). Based on these profiling data, a troubleshooting system can detect subtle anomalies of program behaviors. But, given the sheer volume of profiling data across the stack, analyzing data of all workers at runtime cannot be done at scale. To leverage profiling, engineers must manually reproduce the issues in a small-scale LMT offline, which is time-consuming and not always possible, leading to prolonged recovery time.
We aim to develop a troubleshooting system to effectively localize fine-grained root causes for production issues. The key challenge is to offer fine-grained observability as in offline profiling, while being lightweight for online use cases.
@@
Contributions. We present PerfTracker, the first online performance troubleshooting system that can use fine-grained profiling data for large-scale model training in production. PerfTracker pinpoints root causes of performance issues in both hardware and software components. It scales to modern AI models on large-scale GPU infrastructure— identifying abnormal function executions for a 3,072-GPU LMT only takes 3 minutes (7 minutes for a 1,000,000-GPU LMT). PerfTracker supports all training frameworks using PyTorch. To use PerfTracker, one only needs to import it in the Python source code for LMT without other modifications.
@@
The key insight of PerfTracker is to develop highly efficient differential observability of LMT. PerfTracker does not attempt to directly analyze raw profiling data of all the workers (e.g., compare all function events between all the workers). Instead, PerfTracker summarizes the runtime behavior pattern of causally related LMT functions on the critical path of the observed performance anomaly. We show that it is possible to concisely summarize each function’s runtime behavior patterns using statistical metrics (e.g., execution duration, hardware resource utilization, etc.), without loss of diagnosability. Since the size of runtime behavior patterns is orders of magnitudes smaller than raw profiling data, the abnormal behavior can be efficiently identified by pattern analysis, leveraging the asymmetric architecture of model training.
@@
We have developed PerfTracker as a production service in our LMT infrastructure with $O(10,000)$ GPUs and deployed it for three months (product homepage https://help.aliyun.com/zh/pai/user-guide/perftracker-online-performance-analysis-diagnostic-tool). PerfTracker has been helping our customers to troubleshoot various LMT performance issues which can hardly be addressed by existing practices or state-of-the-art techniques, including hardware and code-level issues.
@@
  • We present PerfTracker as the first profiling-based online performance troubleshooting system for LMT in production.
@@
  • We share our production results of PerfTracker.
Ethics. This work does not raise any ethical issues.
2. Background and Motivation
2.1. Performance Issues in LMT
Large model training (LMT) requires the cooperation of multiple hardware and software modules (Figure 1). The hardware typically includes GPUs (NVI, 2024a, b), CPUs, DRAM, PCIe, NVLinks (NVL, 2024), NICs (CX6, 2024; CX7, 2024), and inter-host networks (Greenberg et al., 2009; Qian et al., 2024). Software modules include GPU operators, communication libraries like NCCL (NCC, 2023), training frameworks (e.g., Megatron (meg, 2019), Nemo (nem, 2024)), containers, etc.. Malfunctions in any of the above components can lead to performance issues, such as throughput decrease, fluctuations, or training blockage. In this paper we do not focus on LMT crashes, which can be explicitly identified by error messages.
Figure 1. The full stack of large model training.
Figure 2. An overview of LMT performance issues.
Figure 2 breaks down LMT performance issues in our production GPU cluster with $O(10,000)$ GPUs in the past nine months. These issues manifest as slow computation and communication, but with diverse root causes—44.4% of them are caused by hardware issues, and 48.2% are software issues.
@@ 2.1. Performance Issues in LMT
Communication slowdowns can stem from physical networks, including NVLink, PCIe, NICs, and switches with optical modules. Although major communication throughput reduction is easy to detect, pinpointing the exact machines or hardware components is much more challenging. The difficulty is because most communications in LMTs are collective operations—a slowdown in any network link can proportionally reduce the communication throughput of many workers simultaneously. Besides, bugs in communication libraries, RDMA, and workload managers (e.g., Slurm (slu, 2024)) can also lead to communication slowdowns.
@@
2.2. Common Practices
Before PerfTracker, we relied on existing tools or techniques to troubleshoot performance issues of LMT (Table 1). However, none of them meets our needs in production.
Monitoring. We deployed online monitors on hardware performance counters of GPUs, PCIe, NVLink, and the inter-host network with second-level sample rates, using existing tools (e.g., DCGM (dcg, 2024), Dynolog (dyn, 2024)) and state-of-the-art techniques (e.g., Minder (Deng et al., 2024) and MegaScale (Jiang et al., 2024)); we also deployed monitors to trace forward and backward time and CUDA execution time. However, as shown in Figure 2, only 29.6% of performance issues can be detected.
The key reason is that online monitors are coarse-grained and are unaware of LMT behavior, creating a fundamental gap between the observation and the root cause. First, performance issues caused by configurations or code are not visible via hardware monitoring. Second, most warnings from monitors do not necessarily indicate performance issues in LMT; many abnormal indicators can also be a result of temporarily high pressure on hardware (e.g., excessive CNPs) or correctable errors (e.g., XID 63,64, and 92), which do not always degrade the end-to-end performance of LMT. Third, even when performance issues are caused by problems covered by our monitors (e.g., hardware issues), they may not be detectable due to coarse-grained online monitoring.
Profiling. Profilers can analyze fine-grained behavior of LMT and are used as a supplement to online monitors. Profilers like Nsight, Torch Profiler, and Python Profiler record all activities in LMT execution with microsecond-level hardware sampling, which provide sufficient information to identify the root cause of sophisticated performance issues.
However, given their significant overhead (e.g., Torch Profiler generates 100+ MB profiling data for one worker per second), profiling can only be done in a small testbed. Engineers must first reproduce the observed performance issues in production in the testbed before using a profiler. However, reproducing production issues is not always possible or can be done timely. Besides problems that only manifest at a large scale, a practical problem is that user code is often confidential and cannot be disclosed to the GPU cluster providers. In practice, engineers often have to reproduce production issues using similar open-source models, which further reduces the chances of reproduction. Although we invested tremendous efforts to reproduce LMT performance issues reported by our customers, there are still 7.4% of performance issues that cannot be reproduced. For the ones that we eventually reproduced, it takes days to months of effort.
Table 1. Performance diagnosis tools for LMT
@@ 2.2. Common Practices
Offline Nsight Systems (nsy, 2024) $<$200kHz $<$1kHz
Torch Profiler (tor, 2024)
Cupti (cup, 2024)
Online DCGM (dcg, 2024) 1Hz
MegaScale (Jiang et al., 2024) $<$1kHz
Dynolog (dyn, 2024) 0.1Hz 0.1kHz
Minder (Deng et al., 2024) $<$1Hz $<$1Hz
Online PerfTracker $<$200kHz $<$1kHz
2.3. Goals and Challenges
Our goal is to significantly improve both efficiency and effectiveness of LMT performance troubleshooting in production at scale, with (1) fine-grained observability: collecting all function execution events during LMT, with fine-granularity hardware sampling, (2) efficiency and scalability: as fine-grained observability leads to large volumes of data, online troubleshooting must be efficient and scalable to large GPU clusters; (3) low overhead: introducing no performance impact on the routine LMT training, and minimizing the requirements for computation power, network bandwidth, and storage. We address the following challenges:
@@ 2.3. Goals and Challenges
    Profiling generates a huge amount of data. The size of profiling data of a single worker, including application-level tracing (e.g., the timeline of function execution) and hardware metrics (e.g., using nsys (nsy, 2024) for GPU, NVLink, and PCIe, in 10k Hz), is typically 100 MB per second. Suppose an LMT requires 10,000 GPUs, a profiler can generate $\sim$1 TB data per second, which is untenable.
@@
    Avoid expensive coordination. In production, hosts are not always perfectly synchronized in clock time. With protocols like Network Time Protocol (NTP) (Mills, 1991), the error could be $\sim$10 milliseconds. Algorithms that compare timestamps of events between workers would not work—a lot of function executions are on the millisecond or even microsecond scale. A decentralized algorithm is needed.
3. Main Idea
The key idea of PerfTracker is to develop highly efficient differential observability of LMT functions, leveraging homogeneous architectures of LMT. We use the term “function” to refer to any procedure in LMT, including Python functions, GPU/CPU kernel functions, memory operations, etc.. Compared to traditional distributed systems, LMT function executions (examples in Appendix A) exhibit two significant characteristics.
@@ 3. Main Idea
    Within a single worker, most low-level functions are executed repeatedly. This is because training involves many identical iterations, and models are typically composed of repetitive submodules (e.g., transformer blocks).
@@
    Runtime behaviors of functions are highly identical across workers, because modern parallelisms (e.g., at data, pipeline, tensor, and expert levels) distribute workloads evenly with frequent synchronization operations.
@@
Ring communication in LMT (like NCCL (NCC, 2023) AllReduce and AllGather) connects workers in ring-like topologies for direct data exchange with only their adjacent neighbors. A ring topology utilizes both NVLinks (intra-host) and GPU-NIC (inter-host) links. In this example, a NCCL AllReduce group with 32 GPUs is distributed on 4 hosts, each hosting 8 GPUs, with every pair of GPUs sharing two bonded NICs. A malfunctioning NIC downgrades a bond by 50% and creates a performance issue for the NCCL AllReduce function. Since multiple rings typically share the intra-host network (e.g., NVLink) but use different NICs for inter-host communication, we focus on each worker’s GPU-NIC throughput during NCCL AllReduce (Figure 3), to distinguish the communication behaviors of different rings and different workers.
@@
    For workers in a ring using the downgraded bond but not directly connecting to this bond (blue nodes in Fig. 4), the average throughput is lower (Fig. 55(b)) with fluctuations.
@@
Figure 3. PCIe bandwidth utilization (from GPU to NIC) during a ring communication without network issues.
Figure 4. The ring topology with and without a problematic link. We illustrate a link with only six workers for simplicity. There are 32 in our experiment.
@@
Figure 5. GPU-NIC throughput pattern. Left figures are the bandwidth utilizations during the execution of a ring communication function, and right figures show enlarged views of local areas for better observation.
Figure 6. Overview of the PerfTracker system.
@@
To explain the fluctuation in Figure 55(b): ring communication always transfers data in small chunks one by one. At each stage, each worker transfers a chunk to the next worker, and does some computation based on the chunk it received. With a slow link, though normal links in the ring can still reach maximal throughput, they have to wait for the slow link after each stage of chunk transmission. So their throughput fluctuates between zero and the maximal, with a high standard deviation. The slow link does not need to wait, so its throughput remains low with a small standard deviation.
Insights. The above example demonstrates three important opportunities: (1) performance issues can be observed by profiling the behavior of function executions; (2) we can troubleshoot performance issues using differential observability that localizes the offending function executions with abnormal behavior (e.g., low average GPU-NIC throughput without fluctuation); and (3) we do not need to analyze fine-grained raw observability data of all the functions; instead, we only need to summarize their runtime behavior patterns. In the ring communication example, each worker only needs to provide two numbers (mean and standard deviation) as the pattern of the ring communication function. In summary:

### v4

# Pengcheng Zhang1, Yang Zhang1, Hanyu Zhao1, Yong Li1, Dennis Cai1, Ennan Zhai1

@@ Abstract
Troubleshooting performance problems of large model training (LMT) is immensely challenging, due to unprecedented scales of modern GPU clusters, the complexity of software-hardware interactions, and the data intensity of the training process. Existing troubleshooting approaches designed for traditional distributed systems or datacenter networks fall short and can hardly apply to real-world training systems. In this paper, we present Eroica, the first online troubleshooting system that provides both fine-grained observation based on profiling, and coverage of all machines in GPU clusters, to diagnose performance issues in production, including both hardware and software problems (or the mixture of both). Eroica effectively summarizes runtime behavior patterns of LMT function executions via online profiling, and leverages differential observability to localize the root cause with minimal production impact. Eroica has been deployed as a production service for large-scale GPU clusters of $\sim$100,000 GPUs for 1.5 years. It has diagnosed a variety of difficult performance issues with 97.5% success.
1 Introduction
@@
Troubleshooting performance problems of Large Model Training (LMT) has become a grand challenge driven by unprecedented increases in AI model sizes and GPU-based training infrastructures. LMT undertakes complex hardware-software interactions; unexpected behavior of hardware components (e.g., GPUs and their interconnects) to software code and configurations (e.g., data loaders, communication, and PyTorch), could lead to performance issues. With rapid scaling of LMT, occurrences of performance problems inevitably increase accordingly, resulting in significant wastage of compute resources and production incidents [jiang2024megascale, HPN, SuperBench].
In practice, performance troubleshooting of production LMT relies on data from either online monitoring or offline profiling. Unfortunately, neither is ideal, and engineers often face difficult trade-offs between observation granularity and cluster-wide coverage. Online monitoring must be coarse-grained to cover the entire GPU cluster, including all LMT training processes (i.e., workers) in all machines, without impacting performance. Offline profiling provides fine-grained observation but incurs substantial overhead and generates a huge amount of data that cannot be consumed in real time, so it cannot be enabled in real large-scale training; instead, engineers usually create a smaller test training environment for profiling, which lacks timeliness and is not always possible.
@@
Most online monitors use vendor tools to collect performance counters of hardware components (e.g., GPUs [dcgm], CPU [pcm], CPU-GPU [dynolog], and NICs [pcm, mstflint, R_PingMesh]) during LMT, or collect timestamps of specific user function executions [ncclprofiler, ebpf]. To avoid production impacts, typically the hardware sampling is only in second granularity, and only a few important user functions can be selected for monitor. However, in our production experience, the coarse-grained observability can only detect general symptoms (e.g., performance degradation in forward/backward phases or in communication modules), but cannot pinpoint fine-grained root causes (e.g., which line of code or which network link caused the degradation).
@@
Offline profiling (e.g., Torch Profiler [torch_profiler], Nsight System [nsys], and Nsight Compute [nsight_compute]) records all function executions (e.g., Python call stacks, CUDA executions) in LMT and high-frequency hardware sampling (e.g., at 200k Hz). Based on these profiling data, a troubleshooting system can detect subtle anomalies in the program behavior of all LMT workers. But, given the sheer volume of profiling data across the stack, analyzing the data of all workers at runtime cannot be done at scale, so it cannot serve as an online troubleshooting approach for production LMTs.
Contributions. We present Eroica, the first online performance troubleshooting system that offers both fine-grained observability as in offline profiling and cluster-wide coverage as in online monitoring (e.g., can be enabled simultaneously at all LMT workers to cover the entire GPU cluster), effectively localizing root causes of LMT performance issues.
@@
Eroica pinpoints root causes of performance issues in both hardware and software components. It scales to modern AI models on large-scale GPU infrastructure— identifying abnormal function executions for a 3,400-GPU LMT only takes 3 minutes (with the ability to analyze a 1,000,000-GPU LMT in 7 minutes). Eroica supports all training frameworks using PyTorch. To use Eroica, one only needs to import it in the Python source code for LMT without other modifications.
@@
The key insight of Eroica is to develop highly efficient differential observability of LMT profiling data. Eroica does not attempt to directly analyze raw profiling data of all workers (e.g., compare all function events between all workers). Instead, Eroica summarizes the runtime behavior pattern of causally related LMT functions on the critical path of the observed performance anomaly. We show that it is possible to concisely summarize each function’s runtime behavior patterns using statistical metrics (e.g., execution duration, hardware resource utilization, etc.), without loss of diagnosability. Since the size of runtime behavior patterns is orders of magnitudes smaller than raw profiling data, the abnormal behavior can be efficiently identified by pattern comparison, leveraging the asymmetric architecture of model training.
@@
We have developed Eroica as a production service in our LMT infrastructure with $\sim$100,000 GPUs and deployed it for 1.5 years. Eroica has been helping our customers to troubleshoot various LMT performance issues which can hardly be addressed by existing practices or state-of-the-art techniques, including hardware and code-level issues.
@@
  • We present Eroica as the first profiling-based online performance troubleshooting system for LMT in production.
@@
  • We share our production results of Eroica.
2 Background and Motivation
2.1  Performance Issues in LMT
Large model training (LMT) requires the cooperation of multiple hardware and software modules (Figure 1). The hardware typically includes GPUs [NVIDIA_H100, NVIDIA_H800], CPUs, DRAM, PCIe, NVLinks [NVLINK], NICs [CX6, CX7], and inter-host networks [greenberg2009vl2, HPN]. Software modules include GPU operators, communication libraries like NCCL [NCCL], training frameworks (e.g., Megatron [megatron], Nemo [nemo]), containers, etc.. Malfunctions in any of the above components can lead to performance issues, such as throughput decrease, fluctuations, or training blockage. In this paper we do not focus on LMT crashes, which can be explicitly identified by error messages.
Figure 2 breaks down all LMT performance issues in our production GPU cluster with $\sim$100,000 GPUs over nine months. These issues manifest as slow computation or communication, but with diverse root causes—44.4% of them are caused by hardware issues, and 48.2% are software issues.
Figure 1: The full stack of large model training.
Figure 2: An overview of LMT performance issues.
@@ 2.1  Performance Issues in LMT
Communication slowdowns can stem from physical networks, including NVLink, PCIe, NICs, and switches with optical modules. Although major communication throughput reduction is easy to detect, pinpointing the exact machines or hardware components is much more challenging. The difficulty is because most communications in LMTs are collective operations—a slowdown in any network link can proportionally reduce the communication throughput of many workers simultaneously. Besides, bugs in communication libraries, RDMA, and workload managers (e.g., Slurm [slurm]) can also lead to communication slowdowns.
@@
2.2  Common Practices
Table 1: Representative performance diagnosis tools for LMT
@@ 2.2  Common Practices
Online DCGM [dcgm] 1Hz
MegaScale [jiang2024megascale] 1kHz
Dynolog* [dynolog] 0.1Hz 0.1kHz
NCCL Profiler [ncclprofiler]
eBPF [ebpf]
Offline Nsight Systems [nsys] 10-200kHz 1kHz
Torch Profiler [torch_profiler]
Online Eroica 10-200kHz 1kHz
*Dynolog supports Torch Profiler as a plugin, to collect Python and kernel information. But, its diagnosis is based only on hardware information and thus is not regarded as a tool based on Python and kernel events.
Before Eroica, we relied on existing tools or techniques to troubleshoot performance issues of LMT (Table 1). However, none of them meets our needs in production.
Monitoring. Our GPU clusters widely deployed online monitors on hardware performance counters of GPUs, PCIe, NVLink, and the inter-host network with second-level sample rates, using existing techniques (e.g., in DCGM [dcgm], Dynolog [dynolog], and MegaScale [jiang2024megascale]). There are also tools that collect function-level information online. NCCL Profiler [ncclprofiler] (and related work such as Mycroft [deng2025mycroft] and Aegis [dong2025evolution]) focuses on collecting collective communication function events, whereas eBPF [ebpf] techniques (e.g., bpftrace [bpftrace]) collect events from system calls or other user-specified functions. However, as shown in Figure 2, only 29.6% of performance issues can be diagnosed and root-caused by these tools.
The key reason is that online monitors are coarse-grained and unaware of detailed behavior, creating a fundamental gap between observations and root causes. First, performance issues caused by configurations or code are invisible to hardware monitoring. Second, most warnings from monitors are false positives [dong2025evolution], they do not necessarily indicate performance issues in LMT; they can also be results of temporarily high pressure on hardware (e.g., excessive CNPs) or correctable errors (e.g., XID 63,64, and 92). Third, some hardware misbehaviors are fine-grained and bursty (e.g., <1ms), which hardware monitors at second-level granularity may miss. Finally, even simple issues, which are supposed to be detected by hardware monitors like link down, still have a chance to evade, because in production, hosts are dynamically added or removed, making it challenging to achieve 100% hardware monitoring coverage [dong2025evolution].
Profiling. Profilers provide fine-grained observation of LMT and are used as a supplement to online monitors. Nsight System and Torch Profiler record all activities in LMT execution with microsecond-level hardware sampling, which provide sufficient information to identify the root cause of sophisticated performance issues.
However, given their significant overhead (e.g., Torch Profiler generates 100+ MB profiling data for one worker per second), profiling can only be done in a small testbed, or for a few specific workers (e.g., only for rank 0) among thousands of workers in an online LMT. Engineers must first reproduce the observed performance issues in production in the testbed before using a profiler. Unfortunately, reproducing production issues is not always possible or timely. Besides problems that only manifest at a large scale, a practical problem is that user code is often confidential and cannot be disclosed to the GPU cluster providers. Engineers often have to reproduce production issues using similar open-source models, which further reduces the chances of reproduction. Although we invested tremendous efforts to reproduce LMT performance issues reported by our customers, there are still 7.4% of performance issues that cannot be reproduced. For the ones that we eventually reproduced, it takes days to months of effort.
2.3  Goals and Challenges
Our goal is to significantly improve both efficiency and effectiveness of LMT performance troubleshooting in production at scale, with (1) fine-grained observability: profiling all function execution events during LMT, with fine-granularity hardware sampling, (2) efficiency and scalability: as fine-grained observability leads to large volumes of profiling data, online troubleshooting must be efficient and scalable to large GPU clusters; (3) low overhead: introducing no performance impact on the routine LMT training, and minimizing the requirements for computation power, network bandwidth, and storage. We address the following challenges:
@@ 2.3  Goals and Challenges
    Profiling generates a huge amount of data. The size of profiling data of a single worker, including application-level tracing (e.g., the timeline of function execution) and hardware metrics (e.g., using nsys [nsys] for GPU, NVLink, and PCIe, in 10k Hz), is typically 100 MB per second. Suppose an LMT requires 10,000 GPUs, a profiler can generate $\sim$1 TB data per second, which is untenable.
@@
    Avoid expensive coordination. In production, hosts are not always perfectly synchronized in clock time. Even with protocols like Network Time Protocol (NTP) [mills1991internet], the error could be $\sim$10 milliseconds. Algorithms that compare timestamps of events between workers would not work—a lot of function executions are on the millisecond or even microsecond scale. A decentralized algorithm is needed.
3 Main Idea
The key idea of Eroica is to develop highly efficient differential observability of LMT functions, leveraging homogeneous architectures of LMT. We use the term “function” to refer to any procedure in LMT, including Python functions, GPU/CPU kernel functions, memory operations, etc.. Compared to traditional systems, LMT function executions exhibit two new characteristics.
@@ 3 Main Idea
    Runtime behaviors of function executions (e.g., the start/end time of execution, and the hardware usage during execution) are expected to be similar across workers or, when not identical (e.g. in heterogeneous computation like expert/pipeline parallelism), to follow a relatively stable distribution, as a model usually has high symmetry to be distributed to multiple workers (examples in Appendix E).
@@
    In our production, once a training job encounters a performance issue, it always persists until remediation. This is because training iterations repeatedly execute the same functions, and neither hardware faults nor code bugs are self-healing, so repeated execution continues to exhibit the same performance degradation.
@@
Ring communication in LMT (like NCCL [NCCL] AllReduce and AllGather) connects workers in ring-like topologies for direct data exchange with only their adjacent neighbors. A ring topology utilizes both NVLinks (intra-host) and GPU-NIC (inter-host) links. In this example, a NCCL AllReduce group with 32 GPUs is distributed on 4 hosts, each hosting 8 GPUs, with every pair of GPUs sharing two bonded NICs. A malfunctioning NIC downgrades a bond by 50% and creates a performance issue for the NCCL AllReduce function. Since multiple rings typically share the intra-host network (e.g., NVLink) but use different NICs for inter-host communication, we focus on each worker’s GPU-NIC throughput during NCCL AllReduce (Figure 3), to distinguish the communication behaviors of different rings and different workers.
@@
    For workers in a ring using the downgraded bond but not directly connecting to this bond (blue nodes in Fig. 4), the average throughput is lower (Figure 55(b)) with fluctuations.
@@
Figure 3: PCIe bandwidth utilization (from GPU to NIC) during a ring communication without network issues.
Figure 4: The ring topology with and without a problematic link. We illustrate a link with only six workers for simplicity. There are 32 in our experiment.
@@
Figure 5: GPU-NIC throughput pattern. Left figures are the bandwidth utilizations during the execution of a ring communication function, and right figures show enlarged views of local areas for better observation.
Figure 6: Overview of the Eroica system.
@@
To explain the fluctuation in Figure 55(b): ring communication always transfers data in small chunks one by one. At each stage, each worker transfers a chunk to the next worker, and does some computation based on the chunk it received. With a slow link in a ring, although other links can still reach maximal throughput, they have to wait for the slow link after each stage of chunk transmission. So their throughput fluctuates between zero and the maximal, with a high standard deviation. The slow link does not need to wait, so its throughput remains low with a small standard deviation as shown in Figure 55(c).
Insights. The above example demonstrates three important opportunities: (1) performance issues can be observed by profiling the behavior of function executions, and we can use differential observability to localize the offending function executions with abnormal behavior (e.g., low average GPU-NIC throughput without fluctuation); and (2) we do not need to analyze fine-grained raw observability data of all the functions; instead, we only need to summarize their runtime behavior patterns. In the ring communication example, each worker only needs to provide two numbers (mean and standard deviation of GPU-NIC throughput) as the pattern of the ring communication function; (3) Fortunately, runtime behavior patterns (e.g., mean and standard deviation) can be independent from absolute timestamps, which can be compared across multiple hosts without any clock time synchronization, solving the Challenge 2 in §2.3. In summary: